A Boeing 747-8i jet needs 1680 m of runway to land. If its loaded mass is 3.33 × 10^5 kg and it lands at a speed of 275 km/hr, what frictional force is needed to stop it?


1
Expert's answer
2021-01-28T19:37:58-0500

We can find the frictional force that is needed to stop the Boeing 747-8i jet, from the work-kinetic energy theorem:


"\Delta KE=KE_f-KE_i=W=-F_{fr}d,""0-\dfrac{1}{2}mv^2=-F_{fr}d,""F_{fr}=-\dfrac{0-\dfrac{1}{2}mv^2}{d},""F_{fr}=-\dfrac{0-\dfrac{1}{2}\cdot3.33\cdot10^5\ kg\cdot(76.38\ \dfrac{m}{s})^2}{1680}=5.78\cdot10^5\ N."